### v1

# On the role of the unitary transformations in Bell inequalities

Abstract
The role of the unitary transformations in the Bell-CHSH inequality is highlighted, in both Quantum Mechanics and relativistic Quantum Field Theory. In the former case, we discuss, through a few examples, how the violation of the Bell-CHSH inequality can be encoded into unitary transformations acting on a given reference set of Bell’s observables. In the latter case, the Bell-CHSH inequality for a real massive scalar field in the vacuum state is considered. After introducing suitable bounded Hermitian operators, we show that using unitary transformations can significantly improve our previous numerical algorithm, leading to an increase in the size of the violation.
I Introduction
The study of Bell inequalities within the contexts of Quantum Mechanics and Quantum Field Theory provides profound insights into the nature of quantum correlations and the interplay between locality, causality, and entanglement. In Quantum Mechanics, Bell inequalities serve as a fundamental test for the nonlocal correlations inherent in entangled states, providing a way to distinguish quantum predictions from those of classical theories rooted in local realism. Extensive theoretical investigations and experimental validations have consistently underscored the nonlocal nature of quantum mechanical systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
Extending the analysis of Bell inequalities to Quantum Field Theory introduces both significant challenges and new perspectives. The combination of Quantum Mechanics and Special Relativity creates a rich and sophisticated framework in which locality and causality form the foundational principles. Here, quantum fields are defined as operator-valued distributions. They interact with a vacuum state whose structure is deeply intricate, as illuminated by the Reeh-Schlieder theorem [11]. This complexity demands the use of advanced mathematical tools, including von Neumann Algebras, the Algebraic Quantum Field Theory and the Tomita-Takesaki modular theory, to rigorously examine entanglement and nonlocality in relativistic quantum systems [12, 13].
In the present work, we aim at pointing out the relevance of the unitary transformations in the study of the violation of a generalization of Bell inequalities, known as the Bell-Clauser-Horne-Shimony-Holt (Bell-CHSH) inequality [14, 1]. For a quantum bipartite system, the Bell-CHSH inequality reads:
$\langle\psi|\;{\cal C}\;|\psi\rangle=\langle\psi|\;(A+A^{\prime})\otimes B+(A-A^{\prime})\otimes B^{\prime}\;|\psi\rangle\;,$ (1)
where $|\psi\rangle$ stands for the state of the system whose Hilbert space is ${\cal H}={\cal H}_{A}\otimes{\cal H}_{B}$. As it is customary, the letters $(A,B)$ refer to Alice and Bob.
A violation of the Bell-CHSH inequality occurs whenever
$2<\Big{|}\langle\psi|\;{\cal C}\;|\psi\rangle\Big{|}\leq 2\sqrt{2}\;,$ (2)
where the bound $2\sqrt{2}$ is the so-called Tsirelson’s bound [15]. The Alice’s and Bob’s operators, $(A,A^{\prime})$ and $(B,B^{\prime})$, are called Bell’s observables and act on ${\cal H}_{A}$ and ${\cal H}_{B}$, respectively. These operators are Hermitian and dichotomic, fulfilling the conditions:
$A$ $=$ $A^{\dagger}\;,\qquad A^{\prime}={A^{\prime}}^{\dagger}\;,\qquad A^{2}={\mathbf{1}}\;,\qquad{A^{\prime}}^{2}={\mathbf{1}}\;,$
$B$ $=$ $B^{\dagger}\;,\qquad B^{\prime}={B^{\prime}}^{\dagger}\;,\qquad B^{2}={\mathbf{1}}\;,\qquad{B^{\prime}}^{2}={\mathbf{1}}\;,$ (3)
and
$[A,B]=[A^{\prime},B]=[A,B^{\prime}]=[A^{\prime},B^{\prime}]=0\;,\qquad[A,A^{\prime}]\neq 0\;,\qquad[B,B^{\prime}]\neq 0\;.$ (4)
The possibility of making use of unitary transformations stems from the observation that the Bell operators are always determined up to unitary transformations, namely
$A$$\rightarrow$ ${\cal U}^{\dagger}_{A}\;A\;{\cal U}_{A}\;,\qquad A^{\prime}\rightarrow{\cal{U^{\prime}}}^{\dagger}_{A}\;A^{\prime}\;{\cal{U^{\prime}}}_{A}\;,$
$B$$\rightarrow$ ${\cal U}^{\dagger}_{B}\;B\;{\cal U}_{B}\;,\qquad B^{\prime}\rightarrow{\cal{U^{\prime}}}^{\dagger}_{B}\;B^{\prime}\;{\cal{U^{\prime}}}_{B}\;,$ (5)
where $({\cal{U}}_{A},{\cal{U^{\prime}}}_{A})$ and $({\cal{U}}_{B},{\cal{U^{\prime}}}_{B})$ are unitaries acting on ${\cal H}_{A}$ and ${\cal H}_{B}$.
As one can easily figure out, these unitary operators might introduce new parameters, which can be adjusted so as to improve the size of the violation. As we shall see, this will be the case of the Bell-CHSH inequality for a relativistic scalar quantum field in the vacuum state.
The paper is organized as follows. In Section (II) we discuss the role of the unitaries in Quantum Mechanics by revisiting a few examples, namely: the spin $1/2$ Bell’s states, the $n00n$ states and the squeezed states. Section (III) is devoted to the Bell-CHSH inequality in Quantum Field Theory. The case of a real massive scalar field in $1+1$ Minkowski spacetime will be scrutinized. After introducing a set of bounded Hermitian operators, the violation of the Bell-CHSH inequality in the vacuum state is established by making use of the Tomita-Takesaki modular theory. Moreover, we show how the use of unitary transformation improves the size of the violation in our numerical setup. Section (IV) collects our conclusion.
II Unitary transformations and the Bell-CHSH inequality in Quantum Mechanics
We begin this section by revisiting the textbook [^1] See [16] for an account on the Bell-CHSH inequality in Quantum Mechanics. example of the entangled spin $1/2$ singlet state
$|\psi_{AB}\rangle=\frac{1}{\sqrt{2}}\left(|+\rangle_{A}\otimes|-\rangle_{B}-|-\rangle_{A}\otimes|+\rangle_{B}\right)\;,$ (6)
which gives maximal violation of the Bell-CHSH inequality:
$\Big{|}\langle\psi_{AB}|\;{\cal C}\;|\psi_{AB}\rangle\Big{|}=2\sqrt{2}\;.$ (7)
The Bell operators are given by
$A$ $=$ $e^{i\alpha}|-\rangle_{A}{}_{A}\langle+|+e^{-i\alpha}|+\rangle_{A}{}_{A}\langle-|\;,\qquad A^{\prime}=e^{i\alpha^{\prime}}|-\rangle_{A}{}_{A}\langle+|+e^{-i\alpha^{\prime}}|+\rangle_{A}{}_{A}\langle-|\;,$
$B$ $=$ $e^{i\beta}|-\rangle_{B}{}_{B}\langle+|+e^{-i\beta}|+\rangle_{B}{}_{B}\langle-|\;,\qquad B^{\prime}=e^{i\beta^{\prime}}|-\rangle_{B}{}_{B}\langle+|+e^{-i\beta^{\prime}}|+\rangle_{B}{}_{B}\langle-|\;,$ (8)
where the parameters $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$ denote the so-called Bell angles. The operators in Eq. (8) fulfill the conditions (3) and (4). A straightforward calculation gives
$\langle\psi_{AB}|\;A\otimes B\;|\psi_{AB}\rangle=-\cos(\alpha-\beta)\;.$ (9)
The usual choice for $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$, corresponding to Eq. (7), is
$\alpha=0\;,\qquad\beta=\frac{\pi}{4}\;,\qquad\alpha^{\prime}=\frac{\pi}{2}\;,\qquad\beta^{\prime}=-\frac{\pi}{4}\;.$ (10)
Looking now at the Bell operators $(A,A^{\prime})$, $(B,B^{\prime})$, one recognizes that they can be obtained from a pair of reference Bell operators, $(A_{0},B_{0})$, upon acting with unitaries. For example, it turns out that
$A={\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,$ (11)
where ${\cal U}_{A}$ stands for the unitary operator
${\cal U}_{A}=e^{-i\alpha|-\rangle_{A}{}_{A}\langle-|}\;,$ (12)
while $A_{0}$ is the Hermitian dichotomic reference operator
$A_{0}=|-\rangle_{A}{}_{A}\langle+|\;+\;|+\rangle_{A}{}_{A}\langle-|\;,\qquad A_{0}=A^{\dagger}_{0}\;,\qquad A_{0}^{2}={\mathbf{1}}\;.$ (13)
In the same vein,
$A^{\prime}$ $=$ $e^{i\alpha^{\prime}|-\rangle_{A}{}_{A}\langle-|}\;A_{0}\;e^{-i\alpha^{\prime}|-\rangle_{A}{}_{A}\langle-|}\;,\qquad B=e^{i\beta|-\rangle_{B}{}_{B}\langle-|}\;B_{0}\;e^{-i\beta|-\rangle_{B}{}_{B}\langle-|}\;,\qquad B^{\prime}=e^{i\beta^{\prime}|-\rangle_{B}{}_{B}\langle-|}\;B_{0}\;e^{-i\beta^{\prime}|-\rangle_{B}{}_{B}\langle-|}\;,$ (14)
with
$B_{0}$ $=$ $|-\rangle_{B}{}_{B}\langle+|\;+\;|+\rangle_{B}{}_{B}\langle-|\;,\qquad B_{0}=B^{\dagger}_{0}\;,\qquad B_{0}^{2}={\mathbf{1}}\;.$ (15)
This example shows that all parameters $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$, from which the violation of the Bell-CHSH inequality originates, can be encoded into unitaries.
As a second example, one might consider the $n00n$ states:
$|\psi_{n}\rangle=\frac{1}{\sqrt{2}}\left(|n_{A}\rangle\otimes|0\rangle_{B}+|0_{A}\rangle\otimes|n\rangle_{B}\right)\;,$ (16)
where $A,B$ correspond to the Hilbert spaces of two harmonic oscillators: ${\cal H}_{A}=\{|n_{A}\rangle=\frac{1}{\sqrt{n!}}({a^{\dagger}})^{n}|0_{A}\rangle\;,\quad n=0,1....\}$ and ${\cal H}_{B}=\{|n_{B}\rangle=\frac{1}{\sqrt{n!}}{({b^{\dagger}})^{n}}|0_{B}\rangle\;,\quad n=0,1....\}$ with
$[a,a^{\dagger}]=1\;,\qquad[b,b^{\dagger}]=1\;,\qquad[a,b]=[a,b^{\dagger}]=0\;.$ (17)
Here, the maximal violation of the Bell-CHSH inequality is achieved by taking as Bell’s observables the following expressions
$A$ $=$ $e^{i\alpha}|n_{A}\rangle\langle 0_{A}|\;+\;e^{-i\alpha}|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;,$
$A^{\prime}$ $=$ $e^{i\alpha^{\prime}}|n_{A}\rangle\langle 0_{A}|\;+\;e^{-i\alpha^{\prime}}|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;,$
$B$ $=$ $e^{i\beta}|n_{B}\rangle\langle 0_{B}|\;+\;e^{-i\beta}|0_{B}\rangle\langle n_{B}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{B}\rangle\langle m_{B}|\;,$
$B^{\prime}$ $=$ $e^{i\beta^{\prime}}|n_{B}\rangle\langle 0_{B}|\;+\;e^{-i\beta^{\prime}}|0_{B}\rangle\langle n_{B}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{B}\rangle\langle m_{B}|\;,$ (18)
where, as usual, $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$ denote the four Bell parameters. The maximal violation follows by noticing that
$\langle\psi_{n}|\;A\otimes B\;|\psi_{n}\rangle=\cos(\alpha+\beta)\;.$ (19)
Moreover, similarly to the case of spin $1/2$, the operators (18) can be cast into the form of Eq. (11), namely
$A={\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}_{A}=e^{-i\alpha N_{A}}\;,\qquad N_{A}=a^{\dagger}a\;,$ (20)
while, for the reference operator, one gets
$A_{0}=|n_{A}\rangle\langle 0_{A}|\;+\;|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;.$ (21)
Analogous expressions can be written down for $(A^{\prime},B,B^{\prime})$. Again, all parameters $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$ are encoded into unitaries.
As a last example, let us consider the squeezed state
$|\lambda\rangle=(1-\lambda^{2})^{1/2}\sum_{n=0}^{\infty}\lambda^{n}|n_{A}n_{B}\rangle\;,\qquad|n_{A}n_{B}\rangle=|n_{A}\rangle\otimes|n_{B}\rangle\;,\qquad\lambda\in[0,1]\;.$ (22)
The Bell’s observables are introduced by making use of the isospin operators [16], i.e. by separating the basis states $\{|n_{A}\rangle,|n_{B}\rangle\}$ into even and odd states:
$|\lambda\rangle=(1-\lambda^{2})^{1/2}\sum_{n=0}^{\infty}\left(\;\lambda^{2n}|{(2n)}_{A}{(2n)}_{B}\rangle\;+\;\lambda^{(2n+1)}|{(2n+1)}_{A}{(2n+1)}_{B}\rangle\right)\;,$ (23)
and by defining
$A$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\alpha}|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;e^{-i\alpha}|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\right)\;,$
$A^{\prime}$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\alpha^{\prime}}|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;e^{-i\alpha^{\prime}}|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\right)\;,$
$B$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\beta}|{(2n+1)}_{B}\rangle\langle(2n)_{B}|\;+\;e^{-i\beta}|{(2n)}_{B}\rangle\langle(2n+1)_{B}|\right)\;,$
$B^{\prime}$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\beta^{\prime}}|{(2n+1)}_{B}\rangle\langle(2n)_{B}|\;+\;e^{-i\beta^{\prime}}|{(2n)}_{B}\rangle\langle(2n+1)_{B}|\right)\;.$ (24)
It turns out that
$\langle\lambda|\;A\otimes B\;|\lambda\rangle=\frac{2\lambda}{1+\lambda^{2}}\cos(\alpha+\beta)\;,$ (25)
yielding maximal violation for $\lambda$ close to 1.
Again
$A$ $=$ ${\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}_{A}=e^{-i\alpha N_{A}}\;,\qquad N_{A}=a^{\dagger}a\;,$
$A_{0}$ $=$ $\sum_{n=0}^{\infty}\left(\;|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\;\right)\;.$ (26)
Other examples may be provided, including entangled coherent states [16] as well as Greenberger–Horne–Zeilinger (GHZ) states [16]. The above considerations show the relevant role that unitaries play in the violation of the Bell-CHSH inequality in Quantum Mechanics.
III Facing the Bell-CHSH in Quantum Field Theory
We are now ready to face the more complex case of the relativistic Quantum Field Theory. We shall rely on the seminal works by S. J. Summers and R. Werner [12, 13, 17], who have established a set of remarkable results concerning the violation of the Bell-CHSH inequality in the vacuum state, see also the recent review [18].
III.1 The real massive scalar field in 1+1 Minkowski spacetime
Let us begin by considering a free massive real scalar field in $1+1$ Minkowski spacetime. In a plane-wave expansion, it reads
$\varphi(t,x)=\int\!\frac{dk}{2\pi}\frac{1}{2\omega_{k}}\left(e^{-ik_{\mu}x^{\mu}}a_{k}+e^{ik_{\mu}x^{\mu}}a^{\dagger}_{k}\right),$ (27)
where $\omega_{k}=k^{0}=\sqrt{k^{2}+m^{2}}$. For the canonical commutation relations, one has
[…]
$[a_{k},a_{q}]$ $=[a^{\dagger}_{k},a^{\dagger}_{q}]=0.$
It is a well established fact that quantum fields have to be treated as operator-valued distributions [11]. As a consequence, they have to be smeared out to yield well-defined operators acting on the Hilbert space, i.e.
$\varphi(h)=\int\!d^{2}x\;\varphi(x)h(x),$ (29)
where $h$ is a real smooth test function with compact support. Using the smeared fields, the Lorentz-invariant inner product is introduced by means of the two-point smeared Wightman function
$\langle f|g\rangle$ $=\langle 0|\varphi(f)\varphi(g)|0\rangle=\frac{i}{2}\Delta_{PJ}(f,g)+H(f,g),$ (30)
where $f$ and $g$ are also real smooth test functions with compact support, and $\Delta_{PJ}(f,g)$ and $H(f,g)$ are the smeared versions of the Pauli-Jordan and Hadamard expressions
$\Delta_{PJ}(f,g)$ $=\int\!d^{2}xd^{2}yf(x)\Delta_{PJ}(x-y)g(y)\;,$
$H(f,g)$ $=\int\!d^{2}xd^{2}yf(x)H(x-y)g(y)\;.$ (31)
Here, $\Delta_{PJ}(x-y)$ and $H(x-y)$ are given by
$\Delta_{PJ}(t,x)$ $=$ $-\frac{1}{2}\;{\rm sign}(t)\;\theta\left(\lambda(t,x)\right)\;J_{0}\left(m\sqrt{\lambda(t,x)}\right)\;,$
$H(t,x)$ $=$ $-\frac{1}{2}\;\theta\left(\lambda(t,x)\right)\;Y_{0}\left(m\sqrt{\lambda(t,x)}\right)+\frac{1}{\pi}\;\theta\left(-\lambda(t,x)\right)\;K_{0}\left(m\sqrt{-\lambda(t,x)}\right)\;,$ (32)
where
$\lambda(t,x)=t^{2}-x^{2}\;,$ (33)
and $(J_{0},Y_{0},K_{0})$ are Bessel functions, while $m$ is the mass parameter.
Both the Hadamard and Pauli-Jordan distributions are Lorentz-invariant. Notably, the Pauli-Jordan distribution $\Delta_{PJ}(x)$ encodes relativistic causality, as it vanishes outside the light cone. Furthermore, $\Delta_{PJ}(x)$ and the Hadamard distribution $H(x)$ exhibit distinct symmetry properties: $\Delta_{PJ}(x)$ is odd under the transformation $x\rightarrow-x$, whereas $H(x)$ is even. When expressed in terms of smeared fields, the commutator of the field operators takes the form $\left[\varphi(f),\varphi(g)\right]=i\Delta_{PJ}(f,g)$. Within this context, the condition $\left[\phi(f),\phi(g)\right]=0,$ whenever the supports of $f$ and $g$ are spacelike separated, encapsulates causality in an elegant way.
III.2 Basics of the Tomita-Takesaki modular theory
The Tomita-Takesaki modular theory serves as a powerful framework for analyzing the Bell-CHSH inequality in Quantum Field Theory [12, 13, 17]. To set the stage, it is worth briefly reviewing some of the basic features of this elegant theoretical structure.
Let ${\cal O}$ stand for an open region of the Minkowski spacetime and let ${\cal M}({\cal O})$ be the space of test functions with support contained in $\cal O$:
${\cal M}({\cal O})=\{f;\;{\rm such\;that}\;supp(f)\subseteq{\cal O}\}.$ (34)
One introduces the symplectic complement of ${\cal M}({\cal O})$ as
${\cal M^{\prime}}({\cal O})=\{g;\;\Delta_{PJ}(g,f)=0\;\;\forall f\in{\cal M}({\cal O})\}.$ (35)
In other words, ${\cal M^{\prime}}({\cal O})$ is given by the set of all test functions for which the smeared Pauli-Jordan expression defined by Eq. (31) vanishes. The usefulness of the symplectic complement ${\cal M^{\prime}}({\cal O})$ relies on the fact that it allows us to rephrase causality as
$\left[\varphi(g),\varphi(f)\right]=0\;,\;\;\forall\;\;g\in{\cal M^{\prime}}({\cal O})\;\;{\rm and}\;\;f\in{\cal M}({\cal O}).$ (36)
The next step is that of introducing a von Neumann algebra ${\cal W}({\cal M})$ of bounded operators supported in ${\cal O}$, equipped with a cyclic and separating vector, which will be addressed in the discussion of von Neumann algebras later. As a concrete example of such an algebra, one may consider the von Neumann algebra generated by the Weyl operators, namely
${\cal W}({\cal M})={\rm span}\;\left\{{\cal A}_{h}\;,supp(h)\in{\cal M}\right\}\;,$ (37)
where ${\cal A}_{h}$ stands for the unitary Weyl operator ${\cal A}_{h}=e^{i{\varphi}(h)}.$ Using the relation $e^{A}\;e^{B}=\;e^{A+B+\frac{1}{2}[A,B]}$, valid for two operators $(A,B)$ commuting with $[A,B]$, one finds that the Weyl operators give rise to the following algebraic structure:
${\cal A}_{h}\;{\cal A}_{h}^{\prime}$ $=$ $e^{-\frac{1}{2}[{\varphi}(h),{\varphi}(h^{\prime})]}\;{\cal A}_{(h+h^{\prime})}=e^{-\frac{i}{2}\Delta_{\textrm{PJ}}(h,h^{\prime})}\;{\cal A}_{(h+h^{\prime})},$
${\cal A}^{\dagger}_{h}$ $=$ ${\cal A}_{(-h)},$ (38)
where $\Delta_{\textrm{PJ}}(h,h^{\prime})$ is the smeared causal Pauli-Jordan expression (31). Setting $\varphi(h)=a_{h}+a^{\dagger}_{h}\;,$ with $(a_{h},a^{\dagger}_{h})$ being the smeared annihilation and creation operators
$a_{h}=\int\!\frac{dk}{2\pi}\frac{1}{2\omega_{k}}h^{*}(\omega_{k},k)a_{k}\;,\qquad a_{h}^{\dagger}=\int\!\frac{dk}{2\pi}\frac{1}{2\omega_{k}}h(\omega_{k},k)a^{\dagger}_{k}\;,\qquad[a_{h},a^{\dagger}_{h^{\prime}}]=\langle h|h^{\prime}\rangle\;,$ (39)
and $h(\omega_{k},k)$ being the Fourier transformation of the test function $h(x)$, it follows that the vacuum expectation value of the operator ${\cal A}_{h}$ turns out to be
$\langle 0|{\cal A}_{h}|0\rangle=\;e^{-\frac{1}{2}{\lVert h\rVert}^{2}},$ (40)
where ${\lVert h\rVert}^{2}\equiv\langle h|h\rangle$ and the vacuum state $|0\rangle$ is defined by $a_{k}|0\rangle=0,\forall k$. In particular, if $supp_{f}$ and $supp_{g}$ are spacelike separated, causality ensures that the Pauli-Jordan function vanishes. Thus, from the above properties, it follows the useful relation
$\langle 0|{\cal A}_{f}{\cal A}_{g}|0\rangle=\langle 0|{\cal A}_{(f+g)}|0\rangle=\;e^{-\frac{1}{2}{\lVert f+g\rVert}^{2}}.$ (41)
A very important property of the von Neumann algebra ${\cal W}({\cal M})$, Eq. (37), generated by the Weyl operators is that, due to the Reeh-Schlieder theorem [11, 19], the vacuum state $|0\rangle$ is both cyclic and separating, meaning that: $i)$ the set of states $\{{\cal A}_{h}|0\rangle$, ${\cal A}_{h}\in{\cal W}({\cal M})\}$ are dense in the Hilbert space; $ii)$ the condition $\{{\cal A}_{h}|0\rangle=0,{\cal A}_{h}\in{\cal W}({\cal M})\}$, implies ${\cal A}_{h}=0$.
In such a situation, one can apply the modular theory of Tomita-Takesaki [12, 13, 17, 18, 19], which will be presented for a generic von Neumann algebra ${\cal W}({\cal M})$ with a cyclic and separating vector state $|\omega\rangle$. To begin, it is helpful to remind the notion of commutant ${\cal W^{\prime}}({\cal M})$ of the von Neumann algebra ${\cal W}({\cal M})$, namely
${\cal W^{\prime}}({\cal M})=\{w^{\prime};\;\;w^{\prime}\;w=w\;w^{\prime}\;\;\;\forall\;w\in{\cal W}({\cal M})\},$ (42)
i.e., ${\cal W^{\prime}}({\cal M})$ contains all elements which commute with each element of ${\cal W}({\cal M})$. Let us also state the so-called Haag’s duality [19, 11, 20], ${\cal W^{\prime}}({\cal M})={\cal W}({\cal M^{\prime}}),$ namely, the commutant ${\cal W^{\prime}}({\cal M})$ coincides with the elements of ${\cal W}({\cal M^{\prime}})$ obtained by taking elements belonging to the symplectic complement ${\cal M^{\prime}}$ of ${\cal M}$. This duality, to our knowledge, has only been proven in the case of free Bose fields [20].
The Tomita-Takesaki construction makes use of an anti-linear unbounded operator $S$ [21] acting on the von Neumann algebra ${\cal W}({\cal M})$ as
$S\;w|\omega\rangle=w^{\dagger}|\omega\rangle,\qquad\forall w\in{\cal W}({\cal M})\;,$ (43)
from which it follows that $S|\omega\rangle=|\omega\rangle$ and $S^{2}=1$. Making use of the polar decomposition [21]
$S=J\;\Delta^{1/2}\;,$ (44)
where $J$ is the anti-linear modular conjugation operator and $\Delta$ is the self-adjoint and positive modular operator, the following properties holds [21]
$J^{\dagger}$ $=$ $J\;,\qquad J^{2}=1\;,$
$J\Delta^{1/2}\;J$ $=$ $\Delta^{-1/2}\;,$
$S^{\dagger}$ $=$ $J\Delta^{-1/2}\;,\qquad S^{\dagger}S^{\dagger}=1\;,$
$\Delta$ $=$ $S^{\dagger}S\;,\qquad\Delta^{-1}=SS^{\dagger}\;.$ (45)
We can now state the renowed Tomita-Takesaki theorem [21], namely:
 $i)$ $J\;{\cal W}({\cal M})\;J={\cal W^{\prime}}({\cal M})$ as well as $J\;{\cal W^{\prime}}({\cal M})\;J={\cal W}({\cal M})$;
 $ii)$ there is a one-parameter family of operators $\Delta^{it}$ with $t\in\mathbb{R}$ which leave ${\cal W}({\cal M})$ invariant, that is:
$\Delta^{it}\;\;{\cal W}({\cal M})\;\Delta^{-it}=\;{\cal W}({\cal M}).$
This theorem has far reaching consequences and finds applications in many areas [22]. As far as the Bell-CHSH inequality is concerned, it provides a powerful way of obtaining Bob’s operators from Alice’s ones by means of the anti-unitary operator $J$. The construction goes as follows: One picks up two test functions $(f,f^{\prime})\in{\cal M}({\cal O})$ and consider the two Alice’s operators $({\cal A}_{f},{\cal A}_{f^{\prime}})$,
${\cal A}_{f}=e^{i{\varphi}(f)}\;,\qquad{\cal A}_{f^{\prime}}=e^{i{\varphi}(f^{\prime})}\;.$ (46)
Thus, for Bob’s operators, we write
$J{\cal A}_{f}J\;,\qquad J{\cal A}_{f^{\prime}}J\;.$ (47)
From the Tomita-Takesaki Theorem, it follows that the two set of operators $({\cal A}_{f},A_{f^{\prime}})$ and $(J{\cal A}_{f}J,J{\cal A}_{f^{\prime}}J)$ fulfill the necessary requirements, since $(J{\cal A}_{f}J,J{\cal A}_{f^{\prime}}J)$ belong to the commutant ${\cal W^{\prime}}({\cal M})$. Also, it is worth underlining that the action of the operators $J$ and $\Delta$ may be lifted directly into the space of the test functions [20, 23, 24], giving rise to an analogue of the Tomita-Takesaki construction, namely,
$J{\cal A}_{f}J=Je^{i{\varphi}(f)}J\equiv e^{-i{\varphi}(jf)}\;,$ (48)
where the test function $jf\in{\cal M^{\prime}}$. Analogously,
$\Delta^{1/2}{\cal A}_{f}\Delta^{-1/2}\equiv e^{i{\varphi}(\delta^{1/2}f)}\;.$ (49)
The operators $(j,\delta)$ are such that [20, 23, 13, 24]
$s$ $=$ $j\delta^{1/2}\;,\qquad s^{2}=1\;,$
$j\delta^{1/2}j$ $=$ $\delta^{-1/2}\;,$
$s^{\dagger}$ $=$ $j\delta^{-1/2}\;,\qquad s^{\dagger}s^{\dagger}=1\;.$ (50)
The operator $j$ is anti-unitary, while $\delta$ is self-adjoint and positive. Moreover, from [13], one learns that, in the case in which the region ${\cal O}$ is a wedge region of the Minkowski spacetime, the spectrum of the operator $\delta$ is the whole positive real line, i.e. $\log(\delta)=\mathbb{R}$. This result follows from the analysis of Bisognano-Wichmann [25] of the Tomita-Takesaki modular operator for wedge regions $({\cal O}_{R},{\cal O}_{L})$
${\cal O}_{R}=\left\{(x,t)\;,x\geq|t|\right\}\;,\qquad{\cal O}_{L}=\left\{(x,t)\;,-x\geq|t|\right\}\;.$ (51)
The wedge regions $({\cal O}_{R},{\cal O}_{L})$ are causal complement of each other and are left invariant by Lorentz boosts. In what follows, we shall always consider the region ${\cal O}$ to be a wedge region. For instance, one may figure out that Alice is located in the right wedge ${\cal O}_{R}$, while Bob is in the left one ${\cal O}_{L}$.
The operators $(s,s^{\dagger})$ have the physical meaning of projecting into the space ${\cal M}$ and its symplectic complement ${\cal M^{\prime}}$, namely, one can show [23, 13, 24] that a test function $f$ belongs to $\cal M$ if and only if
$sf=f\;.$ (52)
Analogously, a test function $g$ belongs to the symplectic complement ${\cal M^{\prime}}$ if and only if
$s^{\dagger}g=g\;.$ (53)
These properties enable us to construct a useful set of test functions for Alice and Bob. Following [12, 13], Alice’s test functions $(f,f^{\prime})$ can be specified as follows: Picking up the spectral subspace of $\delta$ specified by $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]\subset(0,1)$ and introducing the normalized vector $\phi$ belonging to this subspace, one writes
$f=\eta(1+s)\phi\;,\qquad f^{\prime}=\eta^{\prime}(1+s)i\phi\;,$ (54)
where $(\eta,\eta^{\prime})$ are free arbitrary parameters, corresponding to the norms of $(f,f^{\prime})$. According to the setup outlined above, Eq. (54) ensures that $sf=f$ and $sf^{\prime}=f^{\prime}\;.$ Moreover, one checks that $j\phi$ is orthogonal to $\phi$, i.e. $\langle\phi|j\phi\rangle=0$. In fact, from
$\delta^{-1}(j\phi)=j(j\delta^{-1}j)\phi=j(\delta\phi),$ (55)
it follows that the modular conjugation $j$ exchanges the spectral subspace $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$ into $[1/\lambda^{2}-\varepsilon,1/\lambda^{2}+\varepsilon]$, ensuring that $\phi$ and $j\phi$ are orthogonal. Concerning now the pair of Bob’s test functions, they are given by $(jf,jf^{\prime})$. In fact, we have
$s^{\dagger}(jf)=jf\;,\qquad s^{\dagger}(jf^{\prime})=jf^{\prime}\;,$ (56)
meaning that, as required by the relativistic causality, $(jf,jf^{\prime})$ belong to the symplectic complement $\mathcal{M^{\prime}}(\mathcal{O})$.
Finally, taking into account that $\phi$ belongs to the spectral subspace $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$, it follows that [12, 13, 18], in the limit $\varepsilon\rightarrow 0$,
$||f||^{2}$ $=||jf||^{2}=\eta^{2}(1+\lambda^{2})$
$\langle f|jf\rangle$ $=2\eta^{2}\lambda$
$||f^{\prime}||^{2}$ $=||jf^{\prime}||^{2}={\eta^{\prime}}^{2}(1+\lambda^{2})$
$\langle f^{\prime}|jf^{\prime}\rangle$ $=2{\eta^{\prime}}^{2}\lambda$
$\langle f|jf^{\prime}\rangle$ $=0\;.$ (57)
III.3 The violation of the Bell-CHSH inequality in Quantum Field Theory
We have now all tools to discuss the violation of the Bell-CHSH inequality in Quantum Field Theory. Let us proceed with the choice of the Hermitian bounded operators. For such a purpose, we consider the following expressions
$P_{A_{f}}$ $=$ $e^{-i\varphi(f)}\;|0\rangle\langle 0|\;e^{i\varphi(f)}\;,\qquad P_{A_{f^{\prime}}}=e^{-i\varphi(f^{\prime})}\;|0\rangle\langle 0|\;e^{i\varphi(f^{\prime})}\;,$
$P_{B_{jf}}$ $=$ $e^{i\varphi(jf)}\;|0\rangle\langle 0|\;e^{-i\varphi(jf)}\;,\qquad P_{B_{jf^{\prime}}}=e^{i\varphi(jf^{\prime})}\;|0\rangle\langle 0|\;e^{-i\varphi(jf^{\prime})}\;,$ (58)
where $(f,f^{\prime})\in{\cal O}_{R}$ and $(jf,jf^{\prime})\in{\cal O}_{L}$ are Alice’s and Bob’s test functions, respectively. It is easily checked that the above operators are Hermitian. By construction, $(P_{A_{f}},P_{A_{f^{\prime}}})$ commute with $(P_{B_{jf}},P_{B_{jf^{\prime}}})$. Moreover, it turns out that these operators are projectors, a feature which follows from the presence of the vacuum projector |$\ket{0}\bra{0}$ |. In fact,
${P_{A_{f}}}{P_{A_{f}}}$ $=e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}$
$=e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}={P_{A_{f}}}\;.$ (59)
For further use, it is helpful to introduce the correlation function for the projectors of Eq. (58), i.e
$\langle 0|\;{\cal P}\;|0\rangle=\langle 0|\;(P_{A_{f}}+P_{A_{f^{\prime}}})P_{B_{jf}}+(P_{A_{f}}-P_{A_{f^{\prime}}})P_{B_{jf^{\prime}}}\;|0\rangle\;.$ (60)
In particular, by making use of Eqs. (57), it follows that
$\langle 0|\;P_{A_{f}}P_{B_{jf}}\;|0\rangle=e^{-\frac{1}{2}||f||^{2}}\;e^{-\frac{1}{2}||f+jf||^{2}}\;e^{-\frac{1}{2}||jf||^{2}}=e^{-2\eta^{2}((1+\lambda^{2})+\lambda)}\;,$ (61)
so that, for the correlation function (60) one gets
$\langle 0|\;{\cal P}\;|0\rangle=e^{-2\eta^{2}((1+\lambda^{2})+\lambda)}+2e^{-(1+\lambda^{2})(\eta^{2}+{\eta^{\prime}}^{2})}-e^{-2{\eta^{\prime}}^{2}((1+\lambda^{2})+\lambda)}.$ (62)
Remarkably, there is an immediate relationship between projectors and dichotomic operators. If $P$ is a projector, then the operator
$A=\mathbf{1}-2P\;,$ (63)
is dichotomic, namely
$A^{2}=\mathbf{1}+4P^{2}-4P=\mathbf{1}\;.$ (64)
Therefore, for the Hermitian and dichotomic Bell operators, we have
${A_{f}}$ $=$ $\mathbf{1}-2P_{A_{f}}\;,\qquad A_{f^{\prime}}=\mathbf{1}-2P_{A_{f^{\prime}}}\;,$
$B_{jf}$ $=$ $\mathbf{1}-2P_{B_{jf}}\;,\qquad B_{jf^{\prime}}=\mathbf{1}-2P_{B_{jf^{\prime}}}\;,$ (65)
so that, for the Bell-CHSH correlation function in the vacuum state [12, 13, 18] one gets
$\langle 0|\;{\cal C}\;|0\rangle=\langle 0|\;(A_{f}+A_{f^{\prime}})B_{jf}+(A_{f}-A_{f^{\prime}})B_{jf^{\prime}}\;|0\rangle\;.$ (66)
A violation of the Bell-CHSH inequality takes place when
$2<\Big{|}\langle 0|\;{\cal C}\;|0\rangle\Big{|}\leq 2\sqrt{2}\;.$ (67)
Making use of the projectors, it follows that
[…]
This expression yields a maximum violation given by
$\langle 0|\;{\cal C}\;|0\rangle=2.35463\;,$ (69)
corresponding to the parameter values
$\lambda=0.998634\;,\qquad\eta=0.00100492\;,\qquad\eta^{\prime}=0.318599\;.$ (70)
As discussed in the previous Section, we now make use of unitary transformations and introduce new Bell operators $({\hat{A}},{\hat{A}^{\prime}})$ and $({\hat{B}},{\hat{B}^{\prime}})$ defined by
$A_{f}$$\rightarrow$ ${\hat{A}}={\cal U}^{\dagger}_{A}\;A_{f}\;{\cal U}_{A}=e^{-i(\alpha\varphi(f)+\beta\varphi(f^{\prime}))}\;A_{f}\;e^{i(\alpha\varphi(f)+\beta\varphi(f^{\prime}))}\;,$
$A_{f^{\prime}}$$\rightarrow$ ${\hat{A}^{\prime}}={\cal U}^{\dagger}_{A}\;A_{f^{\prime}}\;{\cal U^{\prime}}_{A}=e^{-i(\alpha^{\prime}\varphi(f^{\prime})+\beta^{\prime}\varphi(f))}\;A_{f^{\prime}}\;e^{i(\alpha^{\prime}\varphi(f^{\prime})+\beta^{\prime}\varphi(f))}\;,$
$B_{jf}$$\rightarrow$ ${\hat{B}}={\cal U}^{\dagger}_{B}\;B_{jf}\;{\cal U}_{B}=e^{i(\sigma\varphi(jf)+\tau\varphi(jf^{\prime}))}\;B_{jf}\;e^{-i(\sigma\varphi(jf)+\tau\varphi(jf^{\prime}))}\;,$
$B_{jf^{\prime}}$$\rightarrow$ ${\hat{B}^{\prime}}={\cal U^{\prime}}^{\dagger}_{B}\;B_{jf^{\prime}}\;{\cal U^{\prime}}_{B}=e^{i(\sigma^{\prime}\varphi(jf^{\prime})+\tau^{\prime}\varphi(jf))}\;B_{jf}\;e^{-i(\sigma^{\prime}\varphi(jf^{\prime})+\tau^{\prime}\varphi(jf))}\;,$ (71)
where $(\alpha,\beta,\alpha^{\prime},\beta^{\prime})$, $(\sigma,\tau,\sigma^{\prime},\tau^{\prime})$ are free arbitrary parameters. By construction, the operators $({\hat{A}},{\hat{A}^{\prime}})$ commute with $({\hat{B}},{\hat{B}^{\prime}})$.
It turns out that the introduction of unitary transformations allows for a more efficient scan of the candidates for the Bell operators associated with the highest violation of the Bell-CHSH inequality. This happens because the unitaries increase the parameter space in a consistent way. The end result is that we are able to achieve an increase in the size of the violation with $({\hat{A}},{\hat{A}^{\prime}})$ and $({\hat{B}},{\hat{B}^{\prime}})$ as compared with the corresponding violation obtained with $(A_{f},A_{f^{\prime}})$ and $(B_{jf},B_{jf^{\prime}})$. In fact, a numerical analysis gives us
$\langle 0|\;{\hat{\cal C}}\;|0\rangle=\langle 0|\;({\hat{A}}+{\hat{A}^{\prime}}){\hat{B}}+({\hat{A}}-{\hat{A}^{\prime}}){\hat{B}^{\prime}}\;|0\rangle=2.54066\;,$ (72)
which is attained for
$\lambda$ $=$ $0.999268\;,\qquad\eta=2.71042\;,\qquad\eta^{\prime}=-2.17747\times 10^{-12}\;,\qquad\alpha=-0.977432\;,$
$\beta$ $=$ $-126.675\;,\qquad\sigma=-0.977432\;,\qquad\tau=-126.675\;,\qquad\alpha^{\prime}=-127.182\;,$
$\beta^{\prime}$ $=$ $-0.125863\;,\qquad\sigma^{\prime}=-127.182\;,\qquad\tau^{\prime}=-0.125863\;.$ (73)
The effect of the unitary transformations is well captured by Figs. (1) and (2), in which the behaviours of the correlators $\bra{0}\mathcal{C}\ket{0}$ and $\langle 0|\;{\hat{\cal C}}\;|0\rangle$ as a function of the parameters $(\lambda,\eta)$ are reported.
In Fig. (1), one sees the violation of the Bell-CHSH correlator $\bra{0}\mathcal{C}\ket{0}$ before the use of the unitary transformations. In Fig. (2), the effect of the unitaries is exhibited. Here, the blue surface corresponds to the maximum value attained by $\bra{0}\mathcal{C}\ket{0}$: $2.35463$. The orange surface yields the behaviour of $\bra{0}\hat{\mathcal{C}}\ket{0}$. One sees that the orange surface overcomes the blue one.
Figure 1: Violation of the Bell-CHSH inequality before the use of unitary transformations. The orange surface represents the behaviour of $\bra{0}\mathcal{C}\ket{0}$ as function of the parameters $(\lambda,\eta)$. The blue surface is the classical bound: 2
Figure 2: Behaviour of the Bell-CHSH inequality after the use of unitaries. The orange surface represents the behaviour of $\bra{0}\hat{\mathcal{C}}\ket{0}$ as function of the parameters $(\lambda,\eta)$. The blue surface is related to the maximum value attained by $\bra{0}\mathcal{C}\ket{0}$: $2.35463$.
IV Conclusions
In this work, the role of the unitary transformations has been highlighted. Through a few examples, it has been shown that, in Quantum Mechanics, the Bell operators can be cast into the form
$A={\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}^{\dagger}_{A}{\cal U}_{A}\,=\mathbf{1}=\,{\cal U}_{A}{\cal U}^{\dagger}_{A}\;,\qquad A_{0}=A_{0}^{\dagger}\;,\qquad A_{0}^{2}=\mathbf{1}\;,$ (74)
where $A_{0}$ stands for a reference operator. All parameters leading to the violation of the Bell-CHSH inequality can be encoded in the unitary transformation.
A similar pattern holds in Quantum Field Theory. Here, as exhibited by Eqs. (69) and (72), the size of the violation is considerably improved by making use of unitary transformations, Eqs. (71).
Let us add the following remarks:
  • Looking at Eqs. (71), one is led to figure out that a sequence of unitary transformations could be constructed and applied to the starting operators $(A_{f},A_{f^{\prime}})$ and $(B_{jf},B_{jf^{\prime}})$, Eqs. (58). At each step, the size of the violation should be improved a little bit. Perhaps, this procedure might give a practical way of achieving a final violation close to Tsirelson’s bound: $2\sqrt{2}$.
  • The construction of the bounded Hermitian operators presented in Eqs. (58) looks rather interesting, being easily generalized to a whole family of operators. Let, in fact, $A(\varphi(f))$ denote a generic, non-Hermitian bounded operator in the smeared field $\varphi(f)$. Then, the quantity
    ${\cal A}(\varphi(f))=A^{\dagger}(\varphi(f))\;|0\rangle\langle 0|\;A(\varphi(f))\;,$ (75)
    is Hermitian and bounded.
  • Interestingly enough, the above construction applies to Abelian gauge theories as well. Let $V_{\mu}$ be an Abelian gauge field, subject to the gauge transformation
    $V_{\mu}\rightarrow V_{\mu}+\partial_{\mu}\alpha\;,$ (76)
    and let $V(f)$ be the corresponding smeared field [26]
    $V(f)=\int d^{4}x\;V_{\mu}(x)f^{\mu}(x)\;,$ (77)
    where $\{f^{\mu}\}$ are smooth test functions with compact support. Therefore, the operator
    ${\cal P}_{V}(f)=e^{-iV(f)}\;|0\rangle\langle 0|\;e^{iV(f)}\;,$ (78)
    is gauge invariant, besides of being bounded and Hermitian. Needless to say, expression (78) might open the road for investigating the violation of the Bell-CHSH inequality in the vacuum state of Abelian gauge theories.
Acknowledgments
The authors would like to thank the Brazilian agencies CNPq, CAPES end FAPERJ for financial support. S.P. Sorella, I. Roditi, and M.S. Guimaraes are CNPq researchers under contracts 301030/2019-7, 311876/2021-8, and 309793/2023-8, respectively. F.M. Guedes acknowledges FAPERJ for financial support under the contract SEI-260003/007871/2024.

### v2

@@
-On the role of the unitary transformations in Bell inequalities
+Bell Inequalities and Unitary Dynamics in Quantum Field Theory: Scalar and Proca Fields
 Abstract
-The role of the unitary transformations in the Bell-CHSH inequality is highlighted, in both Quantum Mechanics and relativistic Quantum Field Theory.
-In the former case, we discuss, through a few examples, how the violation of the Bell-CHSH inequality can be encoded into unitary transformations acting on a given reference set of Bell’s observables.
-In the latter case, the Bell-CHSH inequality for a real massive scalar field in the vacuum state is considered.
-After introducing suitable bounded Hermitian operators, we show that using unitary transformations can significantly improve our previous numerical algorithm, leading to an increase in the size of the violation.
+We examine the role of unitary transformations in enhancing violations of the Bell-CHSH inequality in relativistic quantum field theory.
+In quantum systems, all relevant observables can be expressed via unitary rotations of reference operators.
+We extend this perspective to a real massive free scalar field and to the Proca vector field in $(1+1)$ Minkowski spacetime.
+In two dimensions, the Proca theory carries only one physical degree of freedom and is known to be equivalent to the scalar theory, implying that both yield identical Bell-CHSH violations.
+After introducing suitable bounded Hermitian operators, we employ Tomita–Takesaki modular theory to construct spacelike-separated observables in the vacuum.
+We demonstrate that unitary deformations significantly increase the Bell-CHSH violation, approaching the Tsirelson bound, and we report the first explicit evaluation of such violations in a vector-field vacuum.
 I Introduction
-The study of Bell inequalities within the contexts of Quantum Mechanics and Quantum Field Theory provides profound insights into the nature of quantum correlations and the interplay between locality, causality, and entanglement.
-In Quantum Mechanics, Bell inequalities serve as a fundamental test for the nonlocal correlations inherent in entangled states, providing a way to distinguish quantum predictions from those of classical theories rooted in local realism.
-Extensive theoretical investigations and experimental validations have consistently underscored the nonlocal nature of quantum mechanical systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
-Extending the analysis of Bell inequalities to Quantum Field Theory introduces both significant challenges and new perspectives.
-The combination of Quantum Mechanics and Special Relativity creates a rich and sophisticated framework in which locality and causality form the foundational principles.
-Here, quantum fields are defined as operator-valued distributions.
-They interact with a vacuum state whose structure is deeply intricate, as illuminated by the Reeh-Schlieder theorem [11].
-This complexity demands the use of advanced mathematical tools, including von Neumann Algebras, the Algebraic Quantum Field Theory and the Tomita-Takesaki modular theory, to rigorously examine entanglement and nonlocality in relativistic quantum systems [12, 13].
-In the present work, we aim at pointing out the relevance of the unitary transformations in the study of the violation of a generalization of Bell inequalities, known as the Bell-Clauser-Horne-Shimony-Holt (Bell-CHSH) inequality [14, 1].
-For a quantum bipartite system, the Bell-CHSH inequality reads:
+The study of Bell inequalities within the framework of quantum theory has been central to elucidating the profound distinction between quantum correlations and those permitted by local hidden variable theories.
+Originally formulated by Bell in 1964 [1], these inequalities place stringent constraints on statistical correlations predicted by any theory respecting locality and realism.
+Their violation, as predicted by quantum mechanics and experimentally confirmed with increasing sophistication [2, 3, 4, 5, 6, 7, 8, 9, 10, 11], provides compelling evidence for the intrinsically nonlocal character of entanglement.
+While the formulation and experimental verification of Bell inequalities have traditionally been confined to non-relativistic quantum systems with finite degrees of freedom, their extension to the domain of relativistic quantum field theory (QFT) presents both conceptual and technical challenges.
+Unlike finite-dimensional Hilbert space settings, QFT inherently accommodates an infinite number of degrees of freedom and is governed by stringent demands of locality and covariance.
+Fields are operator-valued distributions, and their nontrivial vacuum structure (epitomized by the Reeh-Schlieder theorem [12]) precludes a naive transposition of quantum mechanical results into field-theoretic contexts.
+To rigorously investigate entanglement and nonlocality in relativistic systems, it becomes necessary to employ the mathematical machinery of Algebraic Quantum Field Theory (AQFT), wherein local algebras of observables are associated with spacetime regions.
+Within this formalism, the Tomita–Takesaki modular theory plays a pivotal role, especially in light of Haag duality and the structure of von Neumann algebras generated by local field operators.
+These tools enable a principled construction of observables localized in spacelike separated regions, allowing for the definition of Bell-type inequalities directly in QFT.
+Crucial contributions by Summers and Werner [13, 14, 15] have shown that violations of the Bell-Clauser-Horne-Shimony-Holt (Bell-CHSH) inequality are not only possible in QFT but, remarkably, are generic even in the vacuum state.
+For a non-relativistic quantum bipartite system, the Bell-CHSH inequality reads:
 $\langle\psi|\;{\cal C}\;|\psi\rangle=\langle\psi|\;(A+A^{\prime})\otimes B+(A-A^{\prime})\otimes B^{\prime}\;|\psi\rangle\;,$ (1)
-where $|\psi\rangle$ stands for the state of the system whose Hilbert space is ${\cal H}={\cal H}_{A}\otimes{\cal H}_{B}$. As it is customary, the letters $(A,B)$ refer to Alice and Bob.
+where $|\psi\rangle$ stands for the state of the system whose Hilbert space is ${\cal H}={\cal H}_{A}\otimes{\cal H}_{B}$.
+As it is customary, the letters $(A,B)$ refer to Alice and Bob.
 A violation of the Bell-CHSH inequality occurs whenever
@@
 and
-$[A,B]=[A^{\prime},B]=[A,B^{\prime}]=[A^{\prime},B^{\prime}]=0\;,\qquad[A,A^{\prime}]\neq 0\;,\qquad[B,B^{\prime}]\neq 0\;.$ (4)
+$A,B]=[A^{\prime},B]=[A,B^{\prime}]=[A^{\prime},B^{\prime}]=0\;,\qquad[A,A^{\prime}]\neq 0\;,\qquad[B,B^{\prime}]\neq 0\;.$ (4)
+In the present work, we revisit the Bell-CHSH inequality from the perspective of both quantum mechanics and relativistic QFT, with a particular focus on the role played by unitary transformations.
+In the quantum mechanical setting, we demonstrate through the representative example of the entangled spin-$1/2$ singlet states that the violation of the Bell-CHSH inequality can be encoded in unitary rotations acting on a fixed set of reference observables.
 The possibility of making use of unitary transformations stems from the observation that the Bell operators are always determined up to unitary transformations, namely
@@
 where $({\cal{U}}_{A},{\cal{U^{\prime}}}_{A})$ and $({\cal{U}}_{B},{\cal{U^{\prime}}}_{B})$ are unitaries acting on ${\cal H}_{A}$ and ${\cal H}_{B}$.
-As one can easily figure out, these unitary operators might introduce new parameters, which can be adjusted so as to improve the size of the violation.
-As we shall see, this will be the case of the Bell-CHSH inequality for a relativistic scalar quantum field in the vacuum state.
+These unitary transformations introduce tunable parameters (often corresponding to physical angles or phases) that directly govern the extent of the violation.
+Our central focus, however, lies in extending this analysis to relativistic QFT.
+We first consider the real massive scalar field in $(1+1)$ dimensions, employing the modular framework in Refs.
+[13, 14] to construct bounded Hermitian operators localized in wedge regions.
+Using these operators, we confirm the existence of a Bell-CHSH violation in the vacuum and show that unitary transformations substantially enhance its magnitude, pushing the correlator value closer to Tsirelson’s bound.
+A novel contribution of this paper is the extension of the analysis to the Proca vector field in $(1+1)$ dimensions.
+In higher dimensions, a Proca field propagates massive spin-1 excitations with multiple polarization states.
+However, in two dimensions the transversality condition eliminates any independent polarization, leaving only a single propagating degree of freedom.
+More precisely, the Proca field is dually equivalent to a massive scalar field: one can represent the field strength as $F_{\mu\nu}=m^{2}\,\varepsilon_{\mu\nu}\,\varphi$, or, equivalently, the vector potential as $A^{\mu}=\varepsilon^{\mu\nu}\partial_{\nu}\varphi$, where $\varphi$ is a scalar field and $\varepsilon_{\mu\nu}$ is the two-dimensional Levi-Civita symbol.
+[^1] This duality is exact at the level of free field theories and ensures that the physical Hilbert spaces of the Proca and scalar models are isomorphic.
+The massive scalar field may thus be regarded as the unique dynamical degree of freedom encoded in the Proca field in 1+11+1D.
+By carefully implementing the smearing procedure with transverse test functions and exploiting this equivalence, we show that the Bell-CHSH violation in the Proca vacuum reproduces exactly the scalar case, both before and after the introduction of unitary transformations.
+To the best of our knowledge, this constitutes the first explicit computation of Bell-CHSH violations for a vector field in the QFT vacuum, and it highlights how dualities between field theories can directly shape the manifestation of quantum nonlocality.
 The paper is organized as follows.
-In Section (II) we discuss the role of the unitaries in Quantum Mechanics by revisiting a few examples, namely: the spin $1/2$ Bell’s states, the $n00n$ states and the squeezed states.
-Section (III) is devoted to the Bell-CHSH inequality in Quantum Field Theory.
-The case of a real massive scalar field in $1+1$ Minkowski spacetime will be scrutinized.
-After introducing a set of bounded Hermitian operators, the violation of the Bell-CHSH inequality in the vacuum state is established by making use of the Tomita-Takesaki modular theory.
-Moreover, we show how the use of unitary transformation improves the size of the violation in our numerical setup.
-Section (IV) collects our conclusion.
+In Section II, we review the role of unitary transformations in the Bell-CHSH inequality within quantum mechanics.
+Section III introduces the modular-theoretic construction of localized operators in QFT and applies it to the real scalar field.
+Section IV extends the analysis to the Proca field in $(1+1)$ dimensions.
+Our conclusions and perspectives for future developments are presented in Section V.
 II Unitary transformations and the Bell-CHSH inequality in Quantum Mechanics
-We begin this section by revisiting the textbook [^1] See [16] for an account on the Bell-CHSH inequality in Quantum Mechanics. example of the entangled spin $1/2$ singlet state
+We begin this section by revisiting the textbook [^2] See [17] for an account on the Bell-CHSH inequality in Quantum Mechanics. example of the entangled spin-$1/2$ singlet state
 $|\psi_{AB}\rangle=\frac{1}{\sqrt{2}}\left(|+\rangle_{A}\otimes|-\rangle_{B}-|-\rangle_{A}\otimes|+\rangle_{B}\right)\;,$ (6)
@@
 This example shows that all parameters $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$, from which the violation of the Bell-CHSH inequality originates, can be encoded into unitaries.
-As a second example, one might consider the $n00n$ states:
-$|\psi_{n}\rangle=\frac{1}{\sqrt{2}}\left(|n_{A}\rangle\otimes|0\rangle_{B}+|0_{A}\rangle\otimes|n\rangle_{B}\right)\;,$ (16)
-where $A,B$ correspond to the Hilbert spaces of two harmonic oscillators: ${\cal H}_{A}=\{|n_{A}\rangle=\frac{1}{\sqrt{n!}}({a^{\dagger}})^{n}|0_{A}\rangle\;,\quad n=0,1....\}$ and ${\cal H}_{B}=\{|n_{B}\rangle=\frac{1}{\sqrt{n!}}{({b^{\dagger}})^{n}}|0_{B}\rangle\;,\quad n=0,1....\}$ with
-$[a,a^{\dagger}]=1\;,\qquad[b,b^{\dagger}]=1\;,\qquad[a,b]=[a,b^{\dagger}]=0\;.$ (17)
-Here, the maximal violation of the Bell-CHSH inequality is achieved by taking as Bell’s observables the following expressions
-$A$ $=$ $e^{i\alpha}|n_{A}\rangle\langle 0_{A}|\;+\;e^{-i\alpha}|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;,$
-$A^{\prime}$ $=$ $e^{i\alpha^{\prime}}|n_{A}\rangle\langle 0_{A}|\;+\;e^{-i\alpha^{\prime}}|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;,$
-$B$ $=$ $e^{i\beta}|n_{B}\rangle\langle 0_{B}|\;+\;e^{-i\beta}|0_{B}\rangle\langle n_{B}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{B}\rangle\langle m_{B}|\;,$
-$B^{\prime}$ $=$ $e^{i\beta^{\prime}}|n_{B}\rangle\langle 0_{B}|\;+\;e^{-i\beta^{\prime}}|0_{B}\rangle\langle n_{B}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{B}\rangle\langle m_{B}|\;,$ (18)
-where, as usual, $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$ denote the four Bell parameters. The maximal violation follows by noticing that
-$\langle\psi_{n}|\;A\otimes B\;|\psi_{n}\rangle=\cos(\alpha+\beta)\;.$ (19)
-Moreover, similarly to the case of spin $1/2$, the operators (18) can be cast into the form of Eq. (11), namely
-$A={\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}_{A}=e^{-i\alpha N_{A}}\;,\qquad N_{A}=a^{\dagger}a\;,$ (20)
-while, for the reference operator, one gets
-$A_{0}=|n_{A}\rangle\langle 0_{A}|\;+\;|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;.$ (21)
-Analogous expressions can be written down for $(A^{\prime},B,B^{\prime})$. Again, all parameters $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$ are encoded into unitaries.
-As a last example, let us consider the squeezed state
-$|\lambda\rangle=(1-\lambda^{2})^{1/2}\sum_{n=0}^{\infty}\lambda^{n}|n_{A}n_{B}\rangle\;,\qquad|n_{A}n_{B}\rangle=|n_{A}\rangle\otimes|n_{B}\rangle\;,\qquad\lambda\in[0,1]\;.$ (22)
-The Bell’s observables are introduced by making use of the isospin operators [16], i.e. by separating the basis states $\{|n_{A}\rangle,|n_{B}\rangle\}$ into even and odd states:
-$|\lambda\rangle=(1-\lambda^{2})^{1/2}\sum_{n=0}^{\infty}\left(\;\lambda^{2n}|{(2n)}_{A}{(2n)}_{B}\rangle\;+\;\lambda^{(2n+1)}|{(2n+1)}_{A}{(2n+1)}_{B}\rangle\right)\;,$ (23)
-and by defining
-$A$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\alpha}|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;e^{-i\alpha}|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\right)\;,$
-$A^{\prime}$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\alpha^{\prime}}|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;e^{-i\alpha^{\prime}}|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\right)\;,$
-$B$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\beta}|{(2n+1)}_{B}\rangle\langle(2n)_{B}|\;+\;e^{-i\beta}|{(2n)}_{B}\rangle\langle(2n+1)_{B}|\right)\;,$
-$B^{\prime}$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\beta^{\prime}}|{(2n+1)}_{B}\rangle\langle(2n)_{B}|\;+\;e^{-i\beta^{\prime}}|{(2n)}_{B}\rangle\langle(2n+1)_{B}|\right)\;.$ (24)
-It turns out that
-$\langle\lambda|\;A\otimes B\;|\lambda\rangle=\frac{2\lambda}{1+\lambda^{2}}\cos(\alpha+\beta)\;,$ (25)
-yielding maximal violation for $\lambda$ close to 1.
-Again
-$A$ $=$ ${\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}_{A}=e^{-i\alpha N_{A}}\;,\qquad N_{A}=a^{\dagger}a\;,$
-$A_{0}$ $=$ $\sum_{n=0}^{\infty}\left(\;|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\;\right)\;.$ (26)
-Other examples may be provided, including entangled coherent states [16] as well as Greenberger–Horne–Zeilinger (GHZ) states [16].
+Other examples may be provided, including entangled coherent states, Greenberger-Horne-Zeilinger (GHZ) states, the $N00N$ states and the squeezed state [17, 18].
 The above considerations show the relevant role that unitaries play in the violation of the Bell-CHSH inequality in Quantum Mechanics.
 III Facing the Bell-CHSH in Quantum Field Theory
-We are now ready to face the more complex case of the relativistic Quantum Field Theory.
-We shall rely on the seminal works by S.
+We are now ready to face the more complex case of the relativistic QFT.
+We shall rely on the original works by S.
 J.
@@
 Werner [13, 14, 15], who have established a set of remarkable results concerning the violation of the Bell-CHSH inequality in the vacuum state, see also the recent review [19].
-III.1 The real massive scalar field in 1+1 Minkowski spacetime
-Let us begin by considering a free massive real scalar field in $1+1$ Minkowski spacetime. In a plane-wave expansion, it reads
+III.1 The real massive scalar field in $(1+1)$ Minkowski spacetime
+We consider a free real scalar field of mass $m$ in $(1+1)$-dimensional Minkowski spacetime. Its quantized form, expressed via plane-wave decomposition, reads
 $\varphi(t,x)=\int\!\frac{dk}{2\pi}\frac{1}{2\omega_{k}}\left(e^{-ik_{\mu}x^{\mu}}a_{k}+e^{ik_{\mu}x^{\mu}}a^{\dagger}_{k}\right),$ (16)
-where $\omega_{k}=k^{0}=\sqrt{k^{2}+m^{2}}$. For the canonical commutation relations, one has
-[…]
+where $\omega_{k}=\sqrt{k^{2}+m^{2}}$ and the annihilation and creation operators satisfy the canonical commutation relations
+$[a_{k},a^{\dagger}_{q}]$ $=2\pi\,2\omega_{k}\,\delta(k-q),$ (17)
 $[a_{k},a_{q}]$ $=[a^{\dagger}_{k},a^{\dagger}_{q}]=0.$
-It is a well established fact that quantum fields have to be treated as operator-valued distributions [11].
-As a consequence, they have to be smeared out to yield well-defined operators acting on the Hilbert space, i.e.
-$\varphi(h)=\int\!d^{2}x\;\varphi(x)h(x),$ (29)
-where $h$ is a real smooth test function with compact support. Using the smeared fields, the Lorentz-invariant inner product is introduced by means of the two-point smeared Wightman function
-$\langle f|g\rangle$ $=\langle 0|\varphi(f)\varphi(g)|0\rangle=\frac{i}{2}\Delta_{PJ}(f,g)+H(f,g),$ (30)
-where $f$ and $g$ are also real smooth test functions with compact support, and $\Delta_{PJ}(f,g)$ and $H(f,g)$ are the smeared versions of the Pauli-Jordan and Hadamard expressions
-$\Delta_{PJ}(f,g)$ $=\int\!d^{2}xd^{2}yf(x)\Delta_{PJ}(x-y)g(y)\;,$
-$H(f,g)$ $=\int\!d^{2}xd^{2}yf(x)H(x-y)g(y)\;.$ (31)
-Here, $\Delta_{PJ}(x-y)$ and $H(x-y)$ are given by
-$\Delta_{PJ}(t,x)$ $=$ $-\frac{1}{2}\;{\rm sign}(t)\;\theta\left(\lambda(t,x)\right)\;J_{0}\left(m\sqrt{\lambda(t,x)}\right)\;,$
-$H(t,x)$ $=$ $-\frac{1}{2}\;\theta\left(\lambda(t,x)\right)\;Y_{0}\left(m\sqrt{\lambda(t,x)}\right)+\frac{1}{\pi}\;\theta\left(-\lambda(t,x)\right)\;K_{0}\left(m\sqrt{-\lambda(t,x)}\right)\;,$ (32)
-where
-$\lambda(t,x)=t^{2}-x^{2}\;,$ (33)
-and $(J_{0},Y_{0},K_{0})$ are Bessel functions, while $m$ is the mass parameter.
-Both the Hadamard and Pauli-Jordan distributions are Lorentz-invariant.
-Notably, the Pauli-Jordan distribution $\Delta_{PJ}(x)$ encodes relativistic causality, as it vanishes outside the light cone.
-Furthermore, $\Delta_{PJ}(x)$ and the Hadamard distribution $H(x)$ exhibit distinct symmetry properties: $\Delta_{PJ}(x)$ is odd under the transformation $x\rightarrow-x$, whereas $H(x)$ is even.
-When expressed in terms of smeared fields, the commutator of the field operators takes the form $\left[\varphi(f),\varphi(g)\right]=i\Delta_{PJ}(f,g)$.
-Within this context, the condition $\left[\phi(f),\phi(g)\right]=0,$ whenever the supports of $f$ and $g$ are spacelike separated, encapsulates causality in an elegant way.
+In a rigorous QFT framework, the scalar field is treated as an operator-valued distribution [12], and thus must be smeared with smooth, compactly supported test functions to yield well-defined operators on the Hilbert space.
+Given a test function $h({\bf\it x})\in C_{0}^{\infty}(\mathbb{R}^{1,1})$, with ${\bf\it x}=(t,x)$, the corresponding smeared field operator is defined as
+$\varphi(h)=\int\!d^{2}{\bf\it x}\;\varphi({\bf\it x})h({\bf\it x}).$ (18)
+The Lorentz-invariant vacuum expectation value of two smeared field operators is defined through the two-point Wightman function:
+$\langle f|g\rangle$ $=\langle 0|\varphi(f)\varphi(g)|0\rangle=\frac{i}{2}\Delta_{\rm PJ}(f,g)+H(f,g),$ (19)
+where $f,g\in C_{0}^{\infty}$, and $\Delta_{\text{PJ}}(f,g)$ and $H(f,g)$ denote the smeared Pauli–Jordan and Hadamard distributions, respectively. These are defined by
+$\Delta_{\rm PJ}(f,g)$ $=\int\!d^{2}{\bf\it x}d^{2}{\bf\it y}f({\bf\it x})\Delta_{\rm PJ}({\bf\it x}-{\bf\it y})g({\bf\it y})\;,$
+$H(f,g)$ $=\int\!d^{2}{\bf\it x}d^{2}{\bf\it y}f({\bf\it x})H({\bf\it x}-{\bf\it y})g({\bf\it y})\;.$ (20)
+with distributional kernels given explicitly by
+$\Delta_{\rm PJ}({\bf\it x})$ $=$ $-\frac{1}{2}\;{\rm sign}(t)\;\theta\left(\lambda({\bf\it x})\right)\;J_{0}\left(m\sqrt{\lambda({\bf\it x})}\right)\;,$
+$H({\bf\it x})$ $=$ $-\frac{1}{2}\;\theta\left(\lambda({\bf\it x})\right)\;Y_{0}\left(m\sqrt{\lambda({\bf\it x})}\right)+\frac{1}{\pi}\;\theta\left(-\lambda({\bf\it x})\right)\;K_{0}\left(m\sqrt{-\lambda({\bf\it x})}\right)\;,$ (21)
+where the Lorentz-invariant quantity
+$\lambda({\bf\it x})=t^{2}-x^{2}$ (22)
+determines the causal structure. Here, ($J_{0}$, $Y_{0}$, $K_{0}$) are Bessel functions of the first kind, second kind, and modified second kind, respectively.
+The Pauli-Jordan distribution $\Delta_{\rm PJ}({\bf\it x})$ is Lorentz-invariant and encodes causality, vanishing outside the light cone.
+Moreover, it is odd under ${\bf\it x}\rightarrow-{\bf\it x}$, whereas the Hadamard function $H({\bf\it x})$ is even.
+The commutator of the field operators is thus $\left[\varphi(f),\varphi(g)\right]=i\Delta_{\rm PJ}(f,g)$, ensuring that $\left[\phi(f),\phi(g)\right]=0,$ whenever the supports of $f$ and $g$ are spacelike separated.
+This compactly encodes the principle of microcausality in relativistic field theory.
 III.2 Basics of the Tomita-Takesaki modular theory
-The Tomita-Takesaki modular theory serves as a powerful framework for analyzing the Bell-CHSH inequality in Quantum Field Theory [12, 13, 17].
+The Tomita-Takesaki modular theory serves as a powerful framework for analyzing the Bell-CHSH inequality in QFT [13, 14, 15].
 To set the stage, it is worth briefly reviewing some of the basic features of this elegant theoretical structure.
@@
 One introduces the symplectic complement of ${\cal M}({\cal O})$ as
-${\cal M^{\prime}}({\cal O})=\{g;\;\Delta_{PJ}(g,f)=0\;\;\forall f\in{\cal M}({\cal O})\}.$ (35)
+${\cal M^{\prime}}({\cal O})=\{g;\;\Delta_{\rm PJ}(g,f)=0\;\;\forall f\in{\cal M}({\cal O})\}.$ (24)
 In other words, ${\cal M^{\prime}}({\cal O})$ is given by the set of all test functions for which the smeared Pauli-Jordan expression defined by Eq.
@@
 A very important property of the von Neumann algebra ${\cal W}({\cal M})$, Eq.
-(37), generated by the Weyl operators is that, due to the Reeh-Schlieder theorem [11, 19], the vacuum state $|0\rangle$ is both cyclic and separating, meaning that: $i)$ the set of states $\{{\cal A}_{h}|0\rangle$, ${\cal A}_{h}\in{\cal W}({\cal M})\}$ are dense in the Hilbert space; $ii)$ the condition $\{{\cal A}_{h}|0\rangle=0,{\cal A}_{h}\in{\cal W}({\cal M})\}$, implies ${\cal A}_{h}=0$.
+(26), generated by the Weyl operators is that, due to the Reeh-Schlieder theorem [12, 20], the vacuum state $|0\rangle$ is both cyclic and separating, meaning that: $i)$ the set of states $\{{\cal A}_{h}|0\rangle$, ${\cal A}_{h}\in{\cal W}({\cal M})\}$ are dense in the Hilbert space; ii)ii) the condition $\{{\cal A}_{h}|0\rangle=0,{\cal A}_{h}\in{\cal W}({\cal M})\}$, implies ${\cal A}_{h}=0$.
 In such a situation, one can apply the modular theory of Tomita-Takesaki [13, 14, 15, 19, 20], which will be presented for a generic von Neumann algebra ${\cal W}({\cal M})$ with a cyclic and separating vector state $|\omega\rangle$.
@@
 ${\cal W^{\prime}}({\cal M})=\{w^{\prime};\;\;w^{\prime}\;w=w\;w^{\prime}\;\;\;\forall\;w\in{\cal W}({\cal M})\},$ (31)
-i.e., ${\cal W^{\prime}}({\cal M})$ contains all elements which commute with each element of ${\cal W}({\cal M})$.
-Let us also state the so-called Haag’s duality [19, 11, 20], ${\cal W^{\prime}}({\cal M})={\cal W}({\cal M^{\prime}}),$ namely, the commutant ${\cal W^{\prime}}({\cal M})$ coincides with the elements of ${\cal W}({\cal M^{\prime}})$ obtained by taking elements belonging to the symplectic complement ${\cal M^{\prime}}$ of ${\cal M}$.
-This duality, to our knowledge, has only been proven in the case of free Bose fields [20].
+[…]
 The Tomita-Takesaki construction makes use of an anti-linear unbounded operator $S$ [22] acting on the von Neumann algebra ${\cal W}({\cal M})$ as
@@
 $\Delta$ $=$ $S^{\dagger}S\;,\qquad\Delta^{-1}=SS^{\dagger}\;.$ (34)
-We can now state the renowed Tomita-Takesaki theorem [21], namely:
+We can now state the renowned Tomita-Takesaki theorem [22], namely:
  $i)$ $J\;{\cal W}({\cal M})\;J={\cal W^{\prime}}({\cal M})$ as well as $J\;{\cal W^{\prime}}({\cal M})\;J={\cal W}({\cal M})$;
- $ii)$ there is a one-parameter family of operators $\Delta^{it}$ with $t\in\mathbb{R}$ which leave ${\cal W}({\cal M})$ invariant, that is:
+ ii)ii) there is a one-parameter family of operators $\Delta^{it}$ with $t\in\mathbb{R}$ which leave ${\cal W}({\cal M})$ invariant, that is:
 $\Delta^{it}\;\;{\cal W}({\cal M})\;\Delta^{-it}=\;{\cal W}({\cal M}).$
@@
 For instance, one may figure out that Alice is located in the right wedge ${\cal O}_{R}$, while Bob is in the left one ${\cal O}_{L}$.
-The operators $(s,s^{\dagger})$ have the physical meaning of projecting into the space ${\cal M}$ and its symplectic complement ${\cal M^{\prime}}$, namely, one can show [23, 13, 24] that a test function $f$ belongs to $\cal M$ if and only if
-$sf=f\;.$ (52)
-Analogously, a test function $g$ belongs to the symplectic complement ${\cal M^{\prime}}$ if and only if
+[…]
+sf=f.sf=f\;. (41)
+[…]
 $s^{\dagger}g=g\;.$ (42)
 These properties enable us to construct a useful set of test functions for Alice and Bob.
-Following [12, 13], Alice’s test functions $(f,f^{\prime})$ can be specified as follows: Picking up the spectral subspace of $\delta$ specified by $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]\subset(0,1)$ and introducing the normalized vector $\phi$ belonging to this subspace, one writes
+Following [13, 14], Alice’s test functions $(f,f^{\prime})$ can be specified as follows: Picking up the spectral subspace of $\delta$ specified by $\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]\subset(0,1)$ and introducing the normalized vector $\phi$ belonging to this subspace, one writes
 $f=\eta(1+s)\phi\;,\qquad f^{\prime}=\eta^{\prime}(1+s)i\phi\;,$ (43)
@@
 $\delta^{-1}(j\phi)=j(j\delta^{-1}j)\phi=j(\delta\phi),$ (44)
-it follows that the modular conjugation $j$ exchanges the spectral subspace $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$ into $[1/\lambda^{2}-\varepsilon,1/\lambda^{2}+\varepsilon]$, ensuring that $\phi$ and $j\phi$ are orthogonal.
+it follows that the modular conjugation $j$ exchanges the spectral subspace $\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$ into $1/\lambda^{2}-\varepsilon,1/\lambda^{2}+\varepsilon]$, ensuring that $\phi$ and $j\phi$ are orthogonal.
 Concerning now the pair of Bob’s test functions, they are given by $(jf,jf^{\prime})$.
@@
 meaning that, as required by the relativistic causality, $(jf,jf^{\prime})$ belong to the symplectic complement $\mathcal{M^{\prime}}(\mathcal{O})$.
-Finally, taking into account that $\phi$ belongs to the spectral subspace $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$, it follows that [12, 13, 18], in the limit $\varepsilon\rightarrow 0$,
-$||f||^{2}$ $=||jf||^{2}=\eta^{2}(1+\lambda^{2})$
-$\langle f|jf\rangle$ $=2\eta^{2}\lambda$
-$||f^{\prime}||^{2}$ $=||jf^{\prime}||^{2}={\eta^{\prime}}^{2}(1+\lambda^{2})$
-$\langle f^{\prime}|jf^{\prime}\rangle$ $=2{\eta^{\prime}}^{2}\lambda$
+Finally, taking into account that $\phi$ belongs to the spectral subspace $\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$, it follows that [13, 14, 19], in the limit $\varepsilon\rightarrow 0$,
+[…]
+$\langle f|jf\rangle$ $=2\eta^{2}\lambda,$
+[…]
+$\langle f^{\prime}|jf^{\prime}\rangle$ $=2{\eta^{\prime}}^{2}\lambda,$
 $\langle f|jf^{\prime}\rangle$ $=0\;.$ (46)
 III.3 The violation of the Bell-CHSH inequality in Quantum Field Theory
-We have now all tools to discuss the violation of the Bell-CHSH inequality in Quantum Field Theory.
+We have now all tools to discuss the violation of the Bell-CHSH inequality in QFT.
 Let us proceed with the choice of the Hermitian bounded operators.
@@
 By construction, $(P_{A_{f}},P_{A_{f^{\prime}}})$ commute with $(P_{B_{jf}},P_{B_{jf^{\prime}}})$.
-Moreover, it turns out that these operators are projectors, a feature which follows from the presence of the vacuum projector |$\ket{0}\bra{0}$ |.
+Moreover, it turns out that these operators are projectors, a feature which follows from the presence of the vacuum projector $|\ket{0}\bra{0}$.
 In fact,
-${P_{A_{f}}}{P_{A_{f}}}$ $=e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}$
-$=e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}={P_{A_{f}}}\;.$ (59)
-For further use, it is helpful to introduce the correlation function for the projectors of Eq. (58), i.e
+[…]
+[…]
+It is important to observe that
+$\langle 0|\;\left[P_{A_{f}},P_{B_{jf}}\right]\;|0\rangle$ $=\langle 0|\;\left[P_{A_{f}},P_{B_{jf^{\prime}}}\right]\;|0\rangle=0\;,$
+$\langle 0|\;\left[P_{A_{f^{\prime}}},P_{B_{jf}}\right]\;|0\rangle$ $=\langle 0|\;\left[P_{A_{f^{\prime}}},P_{B_{jf^{\prime}}}\right]\;|0\rangle=0\;,$ (49)
+showing that the operators $(P_{A_{f}},P_{A_{f^{\prime}}})$ commute with $(P_{B_{jf}},P_{B_{jf^{\prime}}})$ in the vacuum state.
+For further use, it is helpful to introduce the correlation function for the projectors of Eq.
+(47), i.e.,
 $\langle 0|\;{\cal P}\;|0\rangle=\langle 0|\;(P_{A_{f}}+P_{A_{f^{\prime}}})P_{B_{jf}}+(P_{A_{f}}-P_{A_{f^{\prime}}})P_{B_{jf^{\prime}}}\;|0\rangle\;.$ (50)
 In particular, by making use of Eqs. (46), it follows that
-$\langle 0|\;P_{A_{f}}P_{B_{jf}}\;|0\rangle=e^{-\frac{1}{2}||f||^{2}}\;e^{-\frac{1}{2}||f+jf||^{2}}\;e^{-\frac{1}{2}||jf||^{2}}=e^{-2\eta^{2}((1+\lambda^{2})+\lambda)}\;,$ (61)
+[…]
 so that, for the correlation function (50) one gets
@@
 $\beta^{\prime}$ $=$ $-0.125863\;,\qquad\sigma^{\prime}=-127.182\;,\qquad\tau^{\prime}=-0.125863\;.$ (63)
-The effect of the unitary transformations is well captured by Figs.
-(1) and (2), in which the behaviours of the correlators $\bra{0}\mathcal{C}\ket{0}$ and $\langle 0|\;{\hat{\cal C}}\;|0\rangle$ as a function of the parameters $(\lambda,\eta)$ are reported.
-In Fig.
-(1), one sees the violation of the Bell-CHSH correlator $\bra{0}\mathcal{C}\ket{0}$ before the use of the unitary transformations.
-In Fig.
-(2), the effect of the unitaries is exhibited.
-Here, the blue surface corresponds to the maximum value attained by $\bra{0}\mathcal{C}\ket{0}$: $2.35463$.
-The orange surface yields the behaviour of $\bra{0}\hat{\mathcal{C}}\ket{0}$.
-One sees that the orange surface overcomes the blue one.
-Figure 1: Violation of the Bell-CHSH inequality before the use of unitary transformations.
-The orange surface represents the behaviour of $\bra{0}\mathcal{C}\ket{0}$ as function of the parameters $(\lambda,\eta)$.
-The blue surface is the classical bound: 2
-Figure 2: Behaviour of the Bell-CHSH inequality after the use of unitaries.
-The orange surface represents the behaviour of $\bra{0}\hat{\mathcal{C}}\ket{0}$ as function of the parameters $(\lambda,\eta)$.
-The blue surface is related to the maximum value attained by $\bra{0}\mathcal{C}\ket{0}$: $2.35463$.
-IV Conclusions
-In this work, the role of the unitary transformations has been highlighted. Through a few examples, it has been shown that, in Quantum Mechanics, the Bell operators can be cast into the form
-$A={\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}^{\dagger}_{A}{\cal U}_{A}\,=\mathbf{1}=\,{\cal U}_{A}{\cal U}^{\dagger}_{A}\;,\qquad A_{0}=A_{0}^{\dagger}\;,\qquad A_{0}^{2}=\mathbf{1}\;,$ (74)
-where $A_{0}$ stands for a reference operator. All parameters leading to the violation of the Bell-CHSH inequality can be encoded in the unitary transformation.
-A similar pattern holds in Quantum Field Theory. Here, as exhibited by Eqs. (69) and (72), the size of the violation is considerably improved by making use of unitary transformations, Eqs. (71).
-Let us add the following remarks:
-  • Looking at Eqs.
-(71), one is led to figure out that a sequence of unitary transformations could be constructed and applied to the starting operators $(A_{f},A_{f^{\prime}})$ and $(B_{jf},B_{jf^{\prime}})$, Eqs.
+[…]
+[…]
+[…]
+Figure 1: Behaviour of the Bell-CHSH inequality before (left panel) and after (rigth panel) the use of unitaries.
+IV The Proca field in ($1+1$)-dimensions
+We now turn to the case of the massive spin-1 Proca field in two-dimensional Minkowski spacetime. The field $V_{\mu}$ is described by the Lagrangian density
+$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}(V)F^{\mu\nu}(V)+\frac{M^{2}}{2}V^{\mu}V_{\mu}\;,$ (64)
+where $F_{\mu\nu}(V)=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}\;$ and the parameter $M$ corresponds to the physical mass of the spin-1 vector field.
+The quantized field admits the plane-wave mode expansion
+$V_{\mu}(t,x)=\int\frac{dk}{2\pi}\frac{1}{2\omega_{k}}[\epsilon_{\mu}(k)a(k)e^{-ik_{\mu}x^{\mu}}+\epsilon_{\mu}(k)a^{\dagger}(k)e^{ik_{\mu}x^{\mu}}]\;,$ (65)
+with dispersion relation $\omega_{k}=\sqrt{k^{2}+M^{2}}$. The polarization vectors satisfy the standard conditions
+$k^{\mu}\,\epsilon_{\mu}(k)=0,\qquad\epsilon^{\mu}(k)\,\epsilon_{\mu}(k)=-1,\qquad\epsilon_{\mu}(k)\,\epsilon_{\nu}(k)=-\left(\eta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{M^{2}}\right)\;,$ (66)
+with $\eta_{\mu\nu}=\textrm{diag}(+,-)$. The creation and annihilation operators fulfill the canonical commutation relations
+$=(2\pi)2\omega_{k}\,\delta(k-q),$ (67)
+$[a_{k},a_{q}]$ $=[a^{\dagger}_{k},a^{\dagger}_{q}]=0\;.$
+A distinctive property of the Proca field is its transversality condition $\partial^{\mu}V_{\mu}=0,$ which implies that only transverse test functions contribute when the field is smeared.
+For a smooth compactly supported function $f_{\mu}({\bf\it x})$, we define
+$V(f)=\int d^{2}{\bf\it x}\,V_{\mu}({\bf\it x})f^{\mu}({\bf\it x})=a_{f}+a^{\dagger}_{f},$ (68)
+with the inner product
+$a_{f},a^{\dagger}_{g}]=\langle f|g\rangle_{V}=-\int\frac{dk}{2\pi}\frac{1}{2\omega_{k}}f^{*\,\mu}(k)g^{\nu}(k)\left(\eta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{M^{2}}\right)\;.$ (69)
+The Fourier transform of the test functions is given by
+$f^{\mu}(k)=\int d^{2}{\bf\it x}\,e^{-ik_{\mu}x^{\mu}}f^{\mu}({\bf\it x})\;.$ (70)
+Equivalently, in configuration space, one finds
+$\langle f|g\rangle_{V}=-\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f^{\mu}({\bf\it x})\left(\eta_{\mu\nu}-\frac{\partial^{{\bf\it x}}_{\mu}\partial^{{\bf\it x}}_{\nu}}{M^{2}}\right)\Delta({\bf\it x}-{\bf\it y})g^{\nu}({\bf\it y})\;,$ (71)
+where
+$\Delta({\bf\it x}-{\bf\it y})=\int\frac{dk}{2\pi}\frac{1}{2\omega_{k}}e^{-ik_{\mu}(x^{\mu}-y^{\mu})}=H({\bf\it x}-{\bf\it y})+\frac{i}{2}\Delta_{\rm PJ}({\bf\it x}-{\bf\it y})\;,$ (72)
+with $H({\bf\it x}-{\bf\it y})$ and $\Delta_{\rm PJ}({\bf\it x}-{\bf\it y})$ denoting the Hadamard and Pauli–Jordan distributions, respectively [cf. Eqs. (19)–(21)].
+To analyze the consequences of transversality, let an arbitrary test function $h_{\mu}({\bf\it x})$ be decomposed into transverse and longitudinal parts,
+$h_{\mu}=h_{\mu}^{T}+h_{\mu}^{L},\qquad h_{\mu}^{T}=\left(\eta_{\mu\nu}-\frac{\partial_{\mu}\partial_{\nu}}{\partial^{2}}\right)h^{\nu},\qquad h_{\mu}^{L}=\frac{\partial_{\mu}\partial_{\nu}}{\partial^{2}}h^{\nu}\;,$ (73)
+where $\partial^{2}=\partial_{0}^{2}-\partial_{1}^{2}.$ Then, the longitudinal component does not contribute to the smeared field:
+$V(h)=\int d^{2}{\bf\it x}\,V^{\mu}h_{\mu}=\int d^{2}{\bf\it x}\,V^{\mu}h^{T}_{\mu}=V(h^{T})\;,$ (74)
+which shows that only transverse test functions are physically relevant. A convenient realization is
+$f_{\mu}(t,x)=\frac{1}{M}(\partial_{1}f,\partial_{0}f)\;,$ (75)
+with $f$ a smooth scalar test function. This ensures that $\partial^{\mu}f_{\mu}=\dfrac{1}{M}(\partial_{0}\partial_{1}f-\partial_{1}\partial_{0}f)=0.$ Substituting into the inner product yields
+$\langle f|g\rangle_{V}$ $=-\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f^{\mu}({\bf\it x})g_{\mu}({\bf\it y})\Delta({\bf\it x}-{\bf\it y})$
+$=-\frac{1}{M^{2}}\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,(\partial_{1}^{{\bf\it x}}f\partial_{1}^{{\bf\it y}}g-\partial_{0}^{{\bf\it x}}f\partial_{0}^{{\bf\it y}}g)\Delta({\bf\it x}-{\bf\it y})$
+$=-\frac{1}{M^{2}}\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f({\bf\it x})g({\bf\it y})(\partial_{1}^{{\bf\it x}}\partial_{1}^{{\bf\it y}}-\partial_{0}^{{\bf\it x}}\partial_{0}^{{\bf\it y}})\Delta({\bf\it x}-{\bf\it y})$
+$=\frac{1}{M^{2}}\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f({\bf\it x})g({\bf\it y})(-\partial_{0}^{2}+\partial_{x}^{2})\Delta({\bf\it x}-{\bf\it y})\;.$ (76)
+Using the Klein–Gordon equation $(\partial_{0}^{2}-\partial_{x}^{2}+M^{2})\Delta({\bf\it x}-{\bf\it y})=0$, this reduces precisely to the scalar-field inner product,
+$\langle f|g\rangle_{V}$ $=\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f({\bf\it x})\Delta({\bf\it x}-{\bf\it y})g({\bf\it y})=\langle f|g\rangle\;,$ (77)
+demonstrating that, in two spacetime dimensions, the massive Proca field is dynamically equivalent to a scalar field.
+The reduction follows from the fact that a vector field in $(1+1)$-dimensions carries a single physical degree of freedom.
+Consequently, the analysis of Bell-CHSH violations parallels the scalar case.
+Introducing the bounded Hermitian projectors
+${\cal P}_{A_{f}}$ $=e^{-iV(f)}\;|0\rangle\langle 0|\;e^{iV(f)}\;,$ (78)
+${\cal P}_{A_{f^{\prime}}}$ $=e^{-iV(f^{\prime})}\;|0\rangle\langle 0|\;e^{iV(f^{\prime})}\;,$
+${\cal P}_{B_{jf}}$ $=e^{iV(jf)}\;|0\rangle\langle 0|\;e^{-iV(jf)}\;,$
+${\cal P}_{B_{jf^{\prime}}}$ $=e^{iV(jf^{\prime})}\;|0\rangle\langle 0|\;e^{-iV(jf^{\prime})}\;,$
+and applying the Tomita-Takesaki modular construction, we obtain for the Bell-CHSH correlator in the vacuum
+[…]
+in complete agreement with the scalar-field result [cf.
+Eq.
+(59)].
+Furthermore, once the additional unitary transformations of Eq.
+(61) are incorporated, the violation is enhanced to the same value reported in Eq.
 (62).
-At each step, the size of the violation should be improved a little bit.
-Perhaps, this procedure might give a practical way of achieving a final violation close to Tsirelson’s bound: $2\sqrt{2}$.
-  • The construction of the bounded Hermitian operators presented in Eqs.
-(58) looks rather interesting, being easily generalized to a whole family of operators.
-Let, in fact, $A(\varphi(f))$ denote a generic, non-Hermitian bounded operator in the smeared field $\varphi(f)$.
-Then, the quantity
-    ${\cal A}(\varphi(f))=A^{\dagger}(\varphi(f))\;|0\rangle\langle 0|\;A(\varphi(f))\;,$ (75)
-    is Hermitian and bounded.
-  • Interestingly enough, the above construction applies to Abelian gauge theories as well. Let $V_{\mu}$ be an Abelian gauge field, subject to the gauge transformation
-    $V_{\mu}\rightarrow V_{\mu}+\partial_{\mu}\alpha\;,$ (76)
-    and let $V(f)$ be the corresponding smeared field [26]
-    $V(f)=\int d^{4}x\;V_{\mu}(x)f^{\mu}(x)\;,$ (77)
-    where $\{f^{\mu}\}$ are smooth test functions with compact support. Therefore, the operator
-    ${\cal P}_{V}(f)=e^{-iV(f)}\;|0\rangle\langle 0|\;e^{iV(f)}\;,$ (78)
-    is gauge invariant, besides of being bounded and Hermitian.
-Needless to say, expression (78) might open the road for investigating the violation of the Bell-CHSH inequality in the vacuum state of Abelian gauge theories.
+Thus, in $(1+1)$-dimensions, the Proca field exhibits no qualitative distinction from the massive scalar field regarding Bell-CHSH violations as expected from their duality relation: both theories possess the same physical content, leading to identical Bell-CHSH violations in the vacuum and exhibiting the same enhancement under unitary deformations.
+V Conclusions
+In this work, we have investigated the role of unitary transformations in optimizing violations of the Bell-CHSH inequality across both quantum-mechanical and field-theoretic settings.
+The central result is that unitary transformations provide a systematic means of enlarging the parameter space of Bell observables, thereby enhancing the extent of nonlocal correlations extracted from a given system.
+At the quantum-mechanical level, we emphasized that Bell observables can always be generated from a fixed reference operator by unitary conjugation, and that the parameters introduced by these unitaries directly control the violation.
+This unifies the treatment of several well-known examples, including singlet states, GHZ states, and coherent-state entanglement, under a common framework.
+Extending this approach to relativistic QFT, we employed the Tomita–Takesaki modular theory to construct localized, bounded Hermitian operators supported in spacelike-separated wedge regions.
+For the real scalar field in $(1+1)$ dimensions, we explicitly computed the Bell-CHSH correlator in the vacuum, recovering a violation of magnitude $\langle 0|\mathcal{C}|0\rangle=2.35463$.
+Introducing unitary deformations raised this value to $\langle 0|\hat{\mathcal{C}}|0\rangle=2.54066$, significantly closer to the Tsirelson bound $2\sqrt{2}$.
+A novel element of our analysis was the inclusion of the Proca field in $(1+1)$ dimensions.
+In this case, the transversality condition restricts the dynamics to a single degree of freedom, and the Proca field admits a dual representation in terms of a massive scalar.
+Concretely, the Proca field in two dimensions admits a Hodge dual representation in terms of the derivative of a scalar, and conversely the scalar can be reconstructed from the dual field tensor.
+This establishes an exact equivalence between the free Proca and scalar theories, ensuring that their Hilbert spaces and correlation structures coincide.
+Our Bell-CHSH analysis makes this correspondence fully explicit: the correlator values obtained for the Proca field reproduce those of the scalar theory, both before and after unitary enhancement.
+This constitutes, to our knowledge, the first demonstration of Bell-CHSH violations in a vector-field vacuum, and it highlights how field-theoretic dualities manifest themselves in the entanglement structure probed by Bell operators.
+Our findings highlight two important messages.
+First, the entanglement structure of the QFT vacuum is sufficiently robust to ensure Bell-CHSH violations across both scalar and vector fields, in line with their duality relations.
+Second, unitary transformations act as a powerful optimization tool, systematically enhancing these violations and providing a route toward approaching Tsirelson’s bound in field-theoretic settings.
+Several directions for future work naturally follow.
+It would be of interest to extend the present construction to higher-dimensional spacetimes, to interacting field theories, and to gauge-invariant sectors where additional constraints arise.
+In particular, Abelian and non-Abelian gauge fields provide fertile ground for testing the universality of our findings.
+A variational approach based on iterated unitary deformations may also provide a systematic method for identifying optimal Bell observables in general QFTs.
+In summary, our results reinforce the deep interplay between algebraic structures, unitarity, and entanglement in quantum field theory.
+They also suggest that unitary freedom remains an underexplored resource for probing the ultimate limits of quantum nonlocality in relativistic systems.
 Acknowledgments
 The authors would like to thank the Brazilian agencies CNPq, CAPES end FAPERJ for financial support.
-S.P.
+S.
+P.
 Sorella, I.
-Roditi, and M.S.
+Roditi, and M.
+S.
 Guimaraes are CNPq researchers under contracts 301030/2019-7, 311876/2021-8, and 309793/2023-8, respectively.
-F.M.
+A.
+F.
+Vieira is supported by a postdoctoral grant from FAPERJ in the Pós-doutorado Nota 10 program, grant No.
+E-26/200.135/2025.
+F.
+M.
 Guedes acknowledges FAPERJ for financial support under the contract SEI-260003/007871/2024.
+D.O.R.
+Azevedo acknowledges PEDECIBA for financial support.